### v1

@@ Abstract
Accurate and efficient climate simulations are crucial for understanding Earth’s evolving climate. However, current general circulation models (GCMs) face challenges in capturing unresolved physical processes, such as cloud and convection. A common solution is to adopt cloud resolving models, that provide more accurate results than the standard subgrid parametrisation schemes typically used in GCMs. However, cloud resolving models, also referred to as super paramtetrizations, remain computationally prohibitive. Hybrid modeling, which integrates deep learning with equation-based GCMs, offers a promising alternative but often struggles with long-term stability and accuracy issues. In this work, we find that water vapor oversaturation during condensation is a key factor compromising the stability of hybrid models. To address this, we introduce CondensNet, a novel neural network architecture that embeds a self-adaptive physical constraint to correct unphysical condensation processes. CondensNet effectively mitigates water vapor oversaturation, enhancing simulation stability while maintaining accuracy and improving computational efficiency compared to super parameterization schemes.
We integrate CondensNet into a GCM to form PCNN-GCM (Physics-Constrained Neural Network GCM), a hybrid deep learning framework designed for long-term stable climate simulations in real-world conditions, including ocean and land. PCNN-GCM represents a significant milestone in hybrid climate modeling, as it shows a novel way to incorporate physical constraints adaptively, paving the way for accurate, lightweight, and stable long-term climate simulations.
@@ 1 Introduction
To address this issue, many attempts focused on enforcing physical constraints within machine learning models, initially on idealized settings beucler2019achieving ; beucler2021enforcing ; yuval2020stable ; yuval2021use . These efforts demonstrated some degree of success, without fully addressing the lack of long-term simulation stability, especially in real-world settings [^1] A real-world setting is a present-day climate model setting, coupled to a land surface model Community Land Model version 4.0 oleson2010technical and forced under prescribed sea surface temperatures and sea ice concentrations hurrell2008new .. More recently, 10-year stable simulations under real land and sea distributions were achieved wang2022stable , although with a trial-and-error approach, and without fully addressing the issue. Recent work has targeted specific aspects, such as relative humidity to prevent excessive moistening fuchs2023torchclim and water condensation, enabling 5-month stable runs behrens2024improving . Physical constraints like global mass and moisture conservation were incorporated into ACE2, a promising DL-based climate emulator with a post-processing correction module watt2024ace2 . Another promising approach used differentiable dynamics to integrate neural networks with numerical methods, improving stability and prediction kochkov2024neural , without directly enforcing physical constraints.
@@
In this study, we show that the regulation of water vapor saturation is critical to obtain long-term stability in hybrid DL climate models, and to avoid oversaturation that in turn leads the simulations to fail. To address this problem, we propose a new neural network architecture, depicted in Figure 1, constituted of two main components: 1) a basic model (BasicNet), learning the cloud representation, and 2) a condensation correction network (ConCorrNet), learning the condensation process. We name this new neural network architecture, composed of the BasicNet and the ConCorrNet,  CondensNet, to emphasize the physical condensation constraint imposed. Notably, unlike simple post‐processing methods that mechanically adjust non‐physical predictions solely to enforce water vapor conservation (often resulting in inaccurate solutions – e.g., significant water vapor biases), our ConCorrNet adaptively refines BasicNet outputs by learning from SPCAM training targets, thereby ensuring that corrections are both physically consistent and accurate, in contrast to those obtained from an unconstrained neural network. With a particular training strategy (detailed in Methods, section 4.1.2), CondensNet can accurately learn cloud physics from super parametrization (CRM) simulations, while imposing a physical constraint that address the water vapor oversaturation issue.
@@
CondensNet together with the chosen host GCM, namely CAM, form the overall hybrid DL-GCM framework that we name PCNN-GCM (Physics Constrained-Neural Network GCM). PCNN-GCM runs under real-world conditions, as it uses Community Land Model version 4.0 lawrence2011parametrization for the land component of the Earth system, and it is forced under prescribed sea surface temperatures and sea ice concentrations according to the Atmospheric Model Intercomparison Project (AMIP) protocol. The new PCNN-GCM framework represents a major step forward in hybrid climate modeling, as it showcases the effective use of adaptive physical constraints to achieve long-term stability without sacrificing accuracy, at a much cheaper computational cost than super parametrization models. It also addresses one of the key drawbacks in previous stable simulations, namely the need for extensive and time-consuming trial-and-error before achieving stability wang2022stable , or the reliance on online training and extensive software engineering resources for developing fully differentiable dynamics kochkov2024neural . We remark that CondensNet can be implemented in any host GCM and can learn any super parametrization (not only SPCAM), hence the naming choice, i.e., PCNN-GCM, of the overall framework.
@@ 2.1 Long-term stability
To understand stability in hybrid climate simulations, we use the same GCM configuration as wang2022stable , referred to as NN-GCM (Neural Network-GCM), and reproduce the stable case, the climate drift case, and the crashed cases. Figure 2, panel I, shows the temporal evolution of total energy, while Figure 2, panel II, shows the temporal evolution of global average total precipitable water content. For both panels, dashed black is the SPCAM reference, gray is CAM5, light green is the stable and unbiased NN-GCM, yellow is NN-GCM with bias, orange is NN-GCM causing climate drift, dark red is an unstable NN-GCM that fails on the 107th simulated day (5105th time step), and dark green is the new PCNN-GCM that produces a stable, unbiased behavior, that follow closely the SPCAM reference. In simulations that fail or produce climate drift, there is an energy surge (Figure 2a), that manifests as an abnormal increase in total precipitable water content (Figure 2d). Indeed, the behavior of the two variables is qualitatively similar.
@@
Figure 2: Total energy (a) and total precipitable water (d) time evolution for different models, including stable (light green), biased (yellow), drifted (orange), and crashed (dark red) ResMLP DL parametrizations, part of NN-GCM, as well as SPCAM (dark green). Total energy (b) and total precipitable water (e) time evolution of Unstable ResMLP models, part of NN-GCM. Total energy (c) and total precipitable water (f) time evolution of stable CondensNet models, part of PCNN-GCM, using as baseline the same configuration of the unstable ResMLP models, to show the effects of CondensNet stabilization properties. Relative humidity for SPCAM reference (g), NN-GCM model failing after 5000 time steps (h), stable NN-GCM, and new PCNN-GCM featuring CondensNet (j).
@@
[…]
If we re-run the unstable NN-GCM simulations shown in Figure 2 (i.e., the crashed NN-GCM simulations depicted in red) using the new PCNN-GCM framework, we now achieve stable simulations without any need for parameter tuning. In particular, we tested PCNN-GCM on six unstable NN-GCM models that led the associated simulations to crash wang2022stable . Figure 2b,e show the total energy and total precipitable water curves before implementing the physical constraint via CondensNet, that is: running the DL parametrization within NN-GCM, which lacks physical constraints. Figure 2c,f show the same curves, using the new PCNN-GCM framework, that imposes a physical constraint on water vapor via CondensNet. The results show that, after employing CondensNet, the total energy curves of the six previously unstable NN-GCM models align closely with SPCAM (ground truth) and remain stable. In addition, the water vapor of PCNN-GCM (Figure 2j) closely resembles the SPCAM reference (Figure 2g), and significantly improves the rather inaccurate results provided by NN-GCM (Figure 2i). We remark that the simulations shown here are under real-world settings that include land and sea components through CLM version 4.0 and AMIP, respectively.
@@ 2.2 Implications on simulation accuracy
We compare PCNN-GCM, against NN-GCM, and CAM5, using SPCAM as the ground truth (and training reference for PCNN-GCM and NN-GCM), where the evaluation period spans from 1 January 1999, to 31 December 2003, whereby the training period ranges from 1 January 1997 to 31 December 1998. CAM5 is introduced as a widely used baseline GCM, and it represents the host GCM adopted by the PCNN-GCM framework. In the results, we refer to CAM5 as GCM=CAM5, to highlight this aspect, and to remark that the CondensNet architecture underlying PCNN-GCM can be seamlessly ported to other host GCMs.
In Table 1, we show the root mean square error (RMSE), calculated as in Equation 11, between each model and the SPCAM reference, for key simulation variables, noting that before calculating the RMSE we calculated the respective means following Equation 9 – see Methods, section 4.2.1 for more details. This comprehensive evaluation quantitatively assesses each model’s ability to reproduce SPCAM’s climate characteristics across multiple spatial and temporal scales, where SPCAM implements a super parametrization that better represents the under-resolved processes than standard subgrid parametrizations kooperman2016robust ; arnold2014effects ; bretherton2014cloud .
@@
Vertical Specific humidity - 0.231 0.144 0.116
@@
The results show how PCNN-GCM produces more accurate results than CAM5 and NN-GCM for all variables related to the water cycle (except for total precipitable water, where the results is close to CAM5, yet significantly better than NN-GCM). In addition, PCNN-GCM is comparable to CAM5 and significantly better than NN-GCM in terms of sensible heat flux, and wind speed (both 10m and vertical). PCNN-GCM performs worse than CAM5 for 2m temperature, yet significantly improving the results of NN-GCM, while it outperforms CAM5 for the vertical distribution of temperature.
@@
The under-performance in 2m temperature predictions stems from the practical difficulty in accessing boundary layer data within the host GCM (i.e., CAM5.2). Since the 2m temperature is diagnostically calculated using surface and lowest-level temperatures, an accurate representation of boundary layer processes (surface-atmosphere interactions, turbulent mixing) is crucial. Currently, both NN-GCM and PCNN-GCM are not accessing planetary boundary layer variables (it would require significant software engineering resources to do so, that is beyond the scope of this work), aspect that affects their capacity to capture near-surface temperature gradients. Yet, we remark that this limitation does not affect the outcomes of this work outcomes – indeed, where boundary-layer information are more easily accessible (e.g., Energy Exascale Earth System Model (E3SM) tang2023fully ), these can be retrieved providing more accurate results for 2m and surface temperature.
@@
We further assessed the performance of PCNN-GCM for two key variables related to the water cycle, given the physical condensation constraints imposed. These are the precipitation field, and the vertical profile of specific humidity in the period 1999–2003, both reported in Figure 3. For the two variables, we present averaged results, obtained as reported in Methods, section 4.2.1, where Figure 3a–d depict precipitation for the reference SPCAM and for CAM5, NN-GCM, and PCNN-GCM, respectively, whereas Figure 3h–k depicts specific humidity. We also show the corresponding differences of each model (i.e., CAM5, NN-GCM, and PCNN-GCM) with respect to SPCAM, for both precipitation (Figure 3e–g) and specific humidity (Figure 3m–o), calculated using Equation 10 in Methods (section 4.2.2), where we also report the RMSE errors in the title of each subfigure. For more details on these calculations, refer to Methods, section 4.2.2.
@@
Figure 3: Precipitation (a–d) and vertical profile of specific humidity (h–k), for SPCAM, CAM5, NN-GCM, and PCNN-GCM, respectively, and corresponding differences with respect to SPCAM reference (e–g, for precipitation, and m–o, for specific humidity). The fields are annual means (1998-2002) computed as reported in Methods 4.2.1, and their differences are computed as in Equation 10, reported in Methods 4.2.2. We also provide error metrics, namely RMSE, for all subfigures related to differences (i.e., e–g for precipitation and m–o for specific humidity).
@@
We present additional results in Supplementary Information A, where we show more detailed analyses of wind speed and temperature (section A.1), precipitation (section A.2), and climate variability (section A.3).

### v2

@@ Abstract
Accurate and efficient climate simulations are crucial for understanding Earth’s evolving climate. However, current general circulation models (GCMs) face challenges in capturing unresolved physical processes, such as cloud and convection. A common solution is to adopt cloud resolving models, that provide more accurate results than the standard subgrid parametrization schemes typically used in GCMs. However, cloud resolving models, also referred to as super parametrizations, remain computationally prohibitive. Hybrid modeling, which integrates deep learning with equation-based GCMs, offers a promising alternative but often struggles with long-term stability and accuracy issues. In this work, we find that water vapor oversaturation during condensation is a key factor compromising the stability of hybrid models. To address this, we introduce CondensNet, a novel neural network architecture that embeds a self-adaptive physical constraint to correct unphysical condensation processes. CondensNet effectively mitigates water vapor oversaturation, enhancing simulation stability while maintaining accuracy and improving computational efficiency compared to super parametrization schemes. We integrate CondensNet into a GCM to form PCNN-GCM (Physics-Constrained Neural Network GCM), a hybrid deep learning framework designed for long-term stable climate simulations in real-world conditions, including ocean and land. PCNN-GCM represents a significant milestone in hybrid climate modeling, as it shows a novel way to incorporate physical constraints adaptively, paving the way for accurate, lightweight, and stable long-term climate simulations.
@@ 1 Introduction
To address this issue, many attempts focused on enforcing physical constraints within machine learning models, initially on idealized settings beucler2019achieving; beucler2021enforcing; yuval2020stable; yuval2021use. These efforts demonstrated some degree of success, without fully addressing the lack of long-term simulation stability, especially in real-world settings (a real-world setting is a present-day climate model setting, coupled to a land surface model Community Land Model version 4.0 oleson2010technical and forced under prescribed sea surface temperatures and sea ice concentrations hurrell2008new). More recently, 10-year stable simulations under real land and sea distributions were achieved wang2022stable, although with a trial-and-error approach, and without fully addressing the issue. Recent work has targeted specific aspects, such as relative humidity to prevent excessive moistening fuchs2023torchclim and water condensation, enabling 5-month stable runs behrens2024improving. Physical constraints like global mass and moisture conservation were incorporated into ACE2, a promising DL-based climate emulator with a post-processing correction module watt2024ace2. Another promising approach used differentiable dynamics to integrate neural networks with numerical methods, improving stability and prediction kochkov2024neural, without directly enforcing physical constraints. Recent results also showed stable simulations for hybrid models with single-module DL parametrizations (i.e., convection-only; radiation-only) balogh2025online; hafner2025stable.
@@
In this study, we show that the regulation of water vapor saturation is critical to obtain long-term stability in hybrid DL climate models, and to avoid oversaturation that in turn leads the simulations to fail. To address this problem, we propose a new neural network architecture, depicted in Figure 1, constituted of two main components: 1) a basic model (BasicNet), learning the cloud representation, and 2) a condensation correction network (ConCorrNet), learning the condensation process. We name this new neural network architecture, composed of the BasicNet and the ConCorrNet,  CondensNet, to emphasize the physical condensation constraint imposed. Here, by cloud representation we refer to the grid-scale parametrization of subgrid cloud and convective processes: the mapping from large-scale prognostic states and forcings to column tendencies at the model resolution. Concretely, in this work it means predicting the water vapor and dry-static-energy tendencies ($\mathrm{d}Q$ and $\mathrm{d}s$) at each vertical level, while radiative effects are produced by a separate DL radiation emulator and are not modified by the condensation constraint. Notably, unlike simple post‐processing methods that mechanically adjust non‐physical predictions solely to enforce water vapor conservation (often resulting in inaccurate solutions – e.g., significant water vapor biases), our ConCorrNet adaptively refines BasicNet outputs by learning from SPCAM training targets, thereby ensuring that corrections are both physically consistent and accurate, in contrast to those obtained from an unconstrained neural network. With a particular training strategy (detailed in Methods, section 4.2), CondensNet can accurately learn cloud physics from super parametrization (CRM) simulations, while imposing a physical constraint that address the water vapor oversaturation issue.
@@
CondensNet together with the chosen host GCM, namely CAM, form the overall hybrid DL-GCM framework that we name PCNN-GCM (Physics Constrained-Neural Network GCM). PCNN-GCM runs under real-world conditions, as it uses Community Land Model version 4.0 lawrence2011parametrization for the land component of the Earth system, and it is forced under prescribed sea surface temperatures and sea ice concentrations according to the Atmospheric Model Intercomparison Project (AMIP) protocol. The new PCNN-GCM framework represents a major step forward in hybrid climate modeling, as it showcases the effective use of adaptive physical constraints to achieve long-term stability without sacrificing accuracy, at a much cheaper computational cost than super parametrization models. PCNN-GCM additionally addresses one of the key drawbacks in previous stable simulations, namely the need for extensive and time-consuming trial-and-error before achieving stability wang2022stable, or the reliance on online training and extensive software engineering resources for developing fully differentiable dynamics kochkov2024neural.
We note that PCNN-GCM is also faster and more accurate, in some respects (such as precipitation and intraseasonal variability), than CAM5, and (unlike CAM) capable of significant GPU acceleration even on a modest number of nodes. For instance, on one 24-core node, PCNN-GCM is nearly 4 times faster than CAM5 when accelerated on GPU – see section 2.3.
We remark that CondensNet is coupled to the CAM/CESM platform; the lightweight Fortran–Python Interface (see Supplementary Information section D) is framework-agnostic and provides a standardized physics–parametrization interface that aligns with how many GCMs exchange model states and tendencies. After model-specific variable and vertical mapping, DL parametrizations can be iterated with minimal host-side engineering, enabling researchers to focus on algorithmic innovation, hence the naming choice, i.e., PCNN-GCM.
@@ 2.1 Long-term stability
To understand stability in hybrid climate simulations, we use the same GCM configuration as wang2022stable, referred to as NN-GCM (Neural Network-GCM), and reproduce the stable case, the climate drift case, and the crashed cases. Figure 2, panel I, shows the temporal evolution of total energy (total energy refers to the vertically integrated atmospheric total energy – internal + kinetic + potential – diagnosed by CAM’s standard energy budget diagnostics, reported in $J\cdot m^{-2}$ and printed in the CESM history/log outputs), while Figure 2, panel II, shows the temporal evolution of global average total precipitable water content. For both panels, dashed black is the SPCAM reference, gray is CAM5, light green is the stable and unbiased NN-GCM, yellow is NN-GCM with bias, orange is NN-GCM causing climate drift, dark red is an unstable NN-GCM that fails on the 107th simulated day (5105th time step), and dark green is the new PCNN-GCM that produces a stable, unbiased behavior, that follow closely the SPCAM reference. In simulations that fail or produce climate drift, there is an energy surge (Figure 2a), that manifests as an abnormal increase in total precipitable water content (Figure 2d). Indeed, the behavior of the two variables is qualitatively similar.
@@
Figure 2: Total energy (a) and total precipitable water (d) time evolution for different models, including stable (light green), biased (yellow), drifted (orange), and crashed (dark red) NN-GCM, PCNN-GCM (dark green), as well as SPCAM (dashed black). Total energy (b) and total precipitable water (e) time evolution of different unstable NN-GCM models. Total energy (c) and total precipitable water (f) time evolution of stable CondensNet models, part of PCNN-GCM, using as baseline the same configuration of the unstable NN-GCM, to show the effects of CondensNet stabilization properties. Relative humidity for SPCAM reference (g), NN-GCM model failing after 5000 time steps (h), stable NN-GCM, and new PCNN-GCM featuring CondensNet (j).
@@
To ensure a physical representation of condensation, we introduce an adaptive physical constraint neural network architecture, termed the condensation correction network (ConCorrNet) – details of ConCorrNet are provided in Methods, section 4. ConCorrNet is tasked to maintain balance in the water cycle process, and it is integrated with a BasicNet i.e., a neural network architecture that is tasked to predict basic tendencies of water vapor (dQQ) and dry-static-energy (dss). ConCorrNet and BasicNet are integrated together to form the new DL parametrization, namely CondensNet, that captures fundamental cloud physics from super parametrization models. CondensNet is then integrated into a GCM, namely CAM version 5.2, to form the overall hybrid DL framework, namely PCNN-GCM. The latter is thoroughly described in Methods, section 4.
If we re-run the unstable NN-GCM simulations shown in Figure 2 (i.e., the crashed NN-GCM simulations depicted in red) using the new PCNN-GCM framework, we now achieve stable simulations without any need for parameter tuning. In particular, we tested PCNN-GCM on six unstable NN-GCM models that led the associated simulations to crash wang2022stable. Figure 2b,e show the total energy and total precipitable water curves before implementing the physical constraint via CondensNet, that is: running the DL parametrization within NN-GCM, which lacks physical constraints. Figure 2c,f show the same curves, using the new PCNN-GCM framework, that imposes a physical constraint on water vapor via CondensNet. The results show that, after employing CondensNet, the total energy curves of the six previously unstable NN-GCM models align closely with SPCAM (ground truth) and remain stable. In addition, the water vapor of PCNN-GCM (Figure 2j) closely resembles the SPCAM reference (Figure 2g), and significantly improves the rather inaccurate results provided by NN-GCM (Figure 2i). In Supplementary Information section A.5, we show 50-year long simulations, to complement the 10-year ones shown in Figure 2. We remark that the simulations shown are under real-world settings that include land and sea components through CLM version 4.0 and AMIP, respectively.
@@ 2.2 Implications on simulation accuracy
We compare PCNN-GCM, against NN-GCM, and CAM5, using SPCAM as the ground truth (and training reference for PCNN-GCM and NN-GCM), where the evaluation period spans from 1 January 1999, to 31 December 2003, whereby the training period ranges from 1 January 1997 to 31 December 1998. CAM5 is introduced as a widely used baseline GCM, and it represents the host GCM adopted by the PCNN-GCM framework. In the results, we refer to CAM5 as GCM=CAM5.
In Table 1, we show the root mean square error (RMSE), calculated as in Equation 12, between each model and the SPCAM reference, for key simulation variables, noting that before calculating the RMSE we calculated the respective means following Equation 10 – see Methods, section 4.3 for more details. This comprehensive evaluation quantitatively assesses each model’s ability to reproduce SPCAM’s climate characteristics across multiple spatial and temporal scales, where SPCAM implements a super parametrization that better represents the under-resolved processes than standard subgrid parametrizations kooperman2016robust; arnold2014effects; bretherton2014cloud.
@@
Zonal-mean specific humidity - 0.231 0.144 0.116
@@
The results show how PCNN-GCM produces more accurate results against SPCAM than CAM5 and NN-GCM for all variables related to the water cycle (except for total precipitable water, where the results is close to CAM5, yet significantly better than NN-GCM). In addition, PCNN-GCM is comparable to CAM5 and significantly better than NN-GCM in terms of sensible heat flux, and wind speed (both 10m and vertical). PCNN-GCM performs worse than CAM5 for 2m temperature, yet significantly improving the results of NN-GCM, while it outperforms CAM5 for the vertical distribution of temperature.
@@
The under-performance in 2m temperature predictions stems from the practical difficulty in accessing boundary layer data within the host GCM (i.e., CAM5.2). Since the 2m temperature is diagnostically calculated using surface and lowest-level temperatures, an accurate representation of boundary layer processes (surface-atmosphere interactions, turbulent mixing) is crucial. Currently, both NN-GCM and PCNN-GCM are not accessing planetary boundary layer variables, such as lowest‑model‑level temperature, surface turbulent fluxes, and a measure of turbulent intensity or mixing, among others. This aspect affects the capacity of NN-GCM and PCNN-GCM to capture near-surface temperature gradients. Accessing these variables would require significant software engineering resources that are beyond the scope of this work. Yet, we remark that this limitation does not affect the outcomes of this study – indeed, where boundary-layer information are more easily accessible (e.g., Energy Exascale Earth System Model (E3SM) tang2023fully), these can be retrieved and can provide a more accurate representation of 2m and surface temperature.
@@
We further assessed the performance of PCNN-GCM for two key variables related to the water cycle, given the physical condensation constraints imposed. These are the precipitation field, and the vertical profile of specific humidity in the period 1999–2003, both reported in Figure 3. For the two variables, we present averaged results, obtained as reported in Methods, section 4.3, where Figure 3a–d depict precipitation for the reference SPCAM and for CAM5, NN-GCM, and PCNN-GCM, respectively, whereas Figure 3h–k depicts specific humidity. We also show the corresponding differences of each model (i.e., CAM5, NN-GCM, and PCNN-GCM) with respect to SPCAM, for both precipitation (Figure 3e–g) and specific humidity (Figure 3m–o), calculated using Equation 11 in Methods (section 4.4), where we also report the RMSE errors in the title of each subfigure. For more details on these calculations, refer to Methods, section 4.4.
@@
Figure 3: Precipitation (a–d) and zonal-mean specific humidity (latitude–pressure; h–k), for SPCAM, CAM5, NN-GCM, and PCNN-GCM, respectively, and corresponding differences with respect to SPCAM reference (e–g, for precipitation, and m–o, for specific humidity). The fields are annual means (1999-2003) computed as reported in Methods 4.3, and their differences are computed as in Equation 11, reported in Methods $4.$ We also provide error metrics, namely RMSE, for all subfigures related to differences (i.e., e–g for precipitation and m–o for specific humidity).
@@
We present additional results in Supplementary Information A, where we show more detailed analyses of wind speed and temperature, precipitation, and climate variability. These results confirm that PCNN-GCM better emulates SPCAM compared to NN-GCM across different features, including: annual means (Supplementary Information section A.1), precipitation distribution (Supplementary Information section A.2), and climate variability (Supplementary Information section A.3 and A.4). We further provide a comparison against ERA5 reanalysis data hersbach2020era5 and TRMM observational data liu2012tropical (Supplementary Information section B). The results of this comparison show SPCAM (and as a consequence PCNN-GCM) producing results relatively different from ERA5 and TRMM. Yet, SPCAM (and PCNN-GCM) shows its strength in capturing tropical wave dynamics as shown by the Wheeler-Kiladis diagrams and the Madden-Julian Oscillation analysis (Supplementary Information sections B.3 and B.4). SPCAM and PCNN-GCM also performs well in terms of precipitation distribution, being closed to ERA5 and TRMM than e.g., CAM5 (Supplementary Information section B.2). Indeed, for both features (tropical wave dynamics and precipitation distribution), SPCAM and PCNN-GCM provide results that are significantly closer to ERA5 and TRMM than CAM5. These results are expected, as the main focus of this work is on obtaining reliably stable hybrid simulations while remaining close to the accuracy of the super parametrization being emulated (i.e., SPCAM), without applying any bias correction for reanalysis and/or observations.